OpenStudy (anonymous):

can some one show me how to do this the radical throws me off. f(x) = sqrt(x-5), g(x) = sqrt(x+4) find (g o f)(x) and give the domain

OpenStudy (anonymous):

g is now a function of f(x), this means that you replace the x in g(x) by f(x) so it looks like this \[\sqrt{5(f(x))}\]

OpenStudy (anonymous):

you know that f(x) = x^2 - 1, so you can substitute that in for f(x) \[\sqrt{5(x^2 - 1)}\]

OpenStudy (anonymous):

and the domain is all values of x that don't result in the number under the radical being negative

OpenStudy (anonymous):

ok i had everything but the parentheses

OpenStudy (anonymous):

yea just remember that x is replaced by EVERYTHING in f(x)

OpenStudy (anonymous):

what do you do if you have 2 radicals?

OpenStudy (anonymous):

that goes in too and then use the rules of exponents to sort it out: \[\sqrt{\sqrt{x}} = (x ^{1/2})^{1/2} = x ^{1/4}\]

OpenStudy (anonymous):

\[=\sqrt[4]{x}\]

OpenStudy (anonymous):

so if f(x)=sqrt{x-5) and g(x)= sqrt(x+4) then f of g would equal the sqrt of the sqrt(x+4)-5?

OpenStudy (anonymous):

I think you got it, the way you wrote it can be interpreted in different ways if you meant \[\sqrt{\sqrt{x+4}-5}\]

OpenStudy (anonymous):

then you got it

OpenStudy (anonymous):

yea i meant that lol

OpenStudy (anonymous):

yay good job :)

OpenStudy (anonymous):

would the answer be the 4th root of x-1?